OpenStudy (anonymous):

Assume that there are 8 different issues of Popular Science, 7 different issues of Time, and 4 different issues of Sports Illustrated, including the December 1st issue, on the rack. You choose 4 of them at random. (1) What is the probability that you choose 3 issues of Popular Science and 1 issue of Time? (2) What is the probability that you choose at least 3 of the Time magazines?

OpenStudy (dumbcow):

P(1) = 4C3 * (7/19)*(8/18)*(7/17)*(6/16)

Directrix (directrix):

P(3 PS and 1 Ti) = [C(8,3)C(7,1)] / C(19,4) = 98/969 = .101 approx

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that didnt work

Directrix (directrix):

See if you wrote the problem correctly. I was expecting a restriction about the Dec 1 SI.

Directrix (directrix):

Mr. Cow has 98/969 = .101 approx

OpenStudy (anonymous):

it never specifies the 1 december 1st issue

Directrix (directrix):

On the first part, Mr. Cow and I agree. Check every word of the problem.

OpenStudy (anonymous):

98/969 is correct .101 is not

OpenStudy (anonymous):

same thing right

Directrix (directrix):

yes, are you to leave answers as fractions?

OpenStudy (anonymous):

i guess so it only worked with 969 but says .101 i tried it multiple times haha

OpenStudy (anonymous):

the next one does specify

OpenStudy (dumbcow):

for next part -> .1174 or 455/3876

OpenStudy (anonymous):

Genius guys

Directrix (directrix):

2) at least 3 T [C(7,3) C(8,1) + C(7,3)C(4,1) +C(7,4)] / C(19,4) = 455 / 3 876 = .117 approx

Directrix (directrix):

You are a genius guy. What's up with my fraction. I'll check.

OpenStudy (anonymous):

thank you so much man you have helped save me